# Goldfish: Monolingual Language Models for 350 Languages

### v1

@@ 6 Conclusion
By design, all of the Goldfish models are monolingual. For low-resource languages, training on closely related languages would likely improve performance (Conneau et al., 2020; Chang et al., 2023). However, adding multilingual data introduces concerns such as the choice of added languages (some languages have more closely related languages in our dataset than others), quantities of added data, and model capacity limitations. To maximize comparability across languages and to allow the models to serve as clearly-defined baselines, we train all Goldfish models monolingually. Of course, language-annotated text datasets inevitably contain mislabeled text, particularly for similar languages (Caswell et al., 2020; Blevins and Zettlemoyer, 2022; Kreutzer et al., 2022). Thus, we cannot guarantee that our models are entirely free from cross-language contamination, although they are monolingual to the best ability of current language identification models.
@@
Because the Goldfish are focused on low-resource languages, we restrict all models to 1GB of training text (after byte premium scaling; Arnett et al., 2024). For the majority of the world’s languages, 1GB is sufficient to include all publicly available text data in the language. At these small dataset sizes, larger models do not appear to provide significant benefit over smaller models (Kaplan et al., 2020; Hoffmann et al., 2022; Chang et al., 2023). Thus, the largest Goldfish model that we train for each language has 125M parameters and is trained on a maximum of 1GB of text. This is the same model size as GPT-1 (Radford et al., 2018) or BERT (Devlin et al., 2019), and the 1GB dataset size is approximately 20% of the dataset size of GPT-1 (Radford et al., 2018).
@@
We evaluate the Goldfish models on FLORES log-perplexity (§4) and three reasoning benchmarks (§5). These are some of the only evaluations that can be used for autoregressive language models in many languages, but they have significant limitations. Perplexity is not necessarily predictive of grammatical text generation (Hu et al., 2020) or complex reasoning capabilities (Levy et al., 2024), but it still provides reasonable signal for model performance (Xia et al., 2023) and it is often used to roughly quantify language model quality (Kaplan et al., 2020; Hoffmann et al., 2022). On the other hand, reasoning benchmarks require annotated datasets and thus often cover fewer languages. One notable exception is Belebele (121 non-English languages; Bandarkar et al., 2024), but even large state-of-the-art models perform quite poorly on Belebele without tuning or few-shot prompting (§5). Thus, our evaluations of model reasoning are not entirely conclusive; we may primarily be measuring heuristics that allow the models to perform only somewhat above chance (arguably, this might still be considered a basic form of “reasoning”). We hope that tractable evaluation datasets with broad language coverage will become increasingly available in the future.

### v2

@@ 6.  Conclusion
By design, all of the Goldfish models are monolingual. For low-resource languages, training on closely related languages would likely improve performance (Conneau et al., 2020b; Chang et al., 2024a). However, adding multilingual data introduces concerns such as the choice of added languages (some languages have more closely related languages in our dataset than others), quantities of added data, and model capacity limitations. To maximize comparability across languages and to allow the models to serve as clearly-defined baselines, we train all Goldfish models monolingually. Of course, language-annotated text datasets inevitably contain mislabeled text, particularly for similar languages (Caswell et al., 2020; Blevins and Zettlemoyer, 2022; Kreutzer et al., 2022). Thus, we cannot guarantee that our models are entirely free from cross-language contamination, although they are monolingual to the best ability of current language identification models.
@@
Because the Goldfish are focused on low-resource languages, we restrict all models to 1GB of training text (after byte premium scaling; Arnett et al., 2024). For the majority of the world’s languages, 1GB is sufficient to include all publicly available text data in the language. At these small dataset sizes, larger models do not appear to provide significant benefit over smaller models (Kaplan et al., 2020; Hoffmann et al., 2022; Chang et al., 2024a). Thus, the largest Goldfish model that we train for each language has 125M parameters and is trained on a maximum of 1GB of text. This is the same model size as GPT-1 (Radford et al., 2018) or BERT (Devlin et al., 2019), and the 1GB dataset size is approximately 20% of the dataset size of GPT-1 (Radford et al., 2018).
@@
We evaluate the Goldfish models on FLORES log-perplexity (§4.1) and four multilingual benchmarks, including three reasoning benchmarks (§4.2). Perplexity is the only evaluation available for autoregressive language models in many languages before instruction-tuning and RLHF, but it has significant limitations. Perplexity is not necessarily predictive of more specific capabilities (Hu et al., 2020a; Levy et al., 2024), although it still provides reasonable signal for model performance (Xia et al., 2023). On the other hand, reasoning benchmarks require annotated datasets and thus often cover fewer languages. Belebele (121 non-English languages; Bandarkar et al., 2024) is an exception, but even state-of-the-art models perform quite poorly on Belebele without instruction-tuning or few-shot prompting (§4.2). Thus, our evaluations of model reasoning are not conclusive; we may primarily be measuring heuristics that allow the models to perform only somewhat above chance (arguably, this might still be considered a basic form of “reasoning”).
Outside of reasoning benchmarks, we also evaluate on MultiBLiMP (Jumelet et al., 2025), which evaluates linguistic knowledge. MultiBLiMP significantly expands the language coverage of multilingual grammaticality benchmarks; however, MultiBLiMP evaluates a very narrow aspect of grammatical knowledge (two types of subject-verb agreement). We hope that tractable evaluation datasets with broad language coverage will become increasingly available in the future to enable more informative evaluation of models in a broader range of languages.